OpenStudy (anonymous):

When playing a game, the winner must roll a number less than 3 or more than 4. What type of events are these? A. Independent B. Dependent C. Mutually exclusive D. Union

OpenStudy (anonymous):

I mean A is my answer

OpenStudy (anonymous):

@pgpilot326 is A correct?

OpenStudy (anonymous):

we must unite them into 1 meaning independent right?

OpenStudy (zarkon):

the event less than 3 and the event more than 4 are not independent

OpenStudy (anonymous):

union?

OpenStudy (zarkon):

try again ;)

OpenStudy (anonymous):

hmmm im gona go with a mautual exclusive?

OpenStudy (anonymous):

they cannot both occur simultaneously.... thus they are _____________ events. know your defs!

OpenStudy (anonymous):

dependent?

OpenStudy (anonymous):

because they depend on eachother?

OpenStudy (anonymous):

you have to work from the definitions. independence and dependence is defined in terms of probabilities. if A and B are independent events, then P(A)*P(B) = P(A and B) if they are dependent, then P(A and B) does not equal P(A)*P(B). mutually exclusive events means that A and B cannot both occur simultaneously. This implies that P(A and B) = 0. Union is just P(A or B) = P(A) + P(B) - P(A and B)

OpenStudy (anonymous):

okay but dependent is the right answer right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay then it has to be mutal exvlusive like i had said above

OpenStudy (anonymous):

event A: player rolls a number less than 3. event B: player rolls more than 4. ask yourself... can A and B occur at the same time... that is, is there some number the player can roll which will satisfy both A and B?

OpenStudy (anonymous):

oh my gosh, it has to be union... I already asked the other person if it was and he said no.

OpenStudy (anonymous):

it's mutually exclusive, silly! you need to learn the definitions, understand what they mean and then learn to apply them. You can do it!